OpenStudy (anonymous):

The function h(x) has y-intercept 3 and goes through the point (a,a+3). what is the formula?

OpenStudy (campbell_st):

well all you need to do is find the slope as you know 2 points, one of which is an intercept the points are (0,3) and (a, a + 3) so the slope is m = (a + 3 - 3)/(a - 0) then you can substitute into the equation y = mx + b where m = slope of the line and b is the y-intercept

OpenStudy (anonymous):

But this is an exponential problem

OpenStudy (campbell_st):

where does it say its an exponential equation... ?

OpenStudy (campbell_st):

ok... so the then for the exponential equation when x = 0 y = 3 as the intercept so the form is y = 3e^kx